# MVG-CRPS: A Robust Loss Function for Multivariate Probabilistic Forecasting

### v1

@@ Abstract
In probabilistic time series forecasting, the multivariate Gaussian (MVG) distribution is widely used as predictive distribution for correlated continuous random variables. Current deep probabilistic models typically employ neural networks to parameterize the mean vector and covariance matrix of the distribution, with log-score (i.e., negative log-likelihood) as the default loss function. However, log-score is highly sensitive to outliers, leading to significant errors when anomalies are present in the data. Motivated by the use of the continuous ranked probability score (CRPS) in learning univariate distributions, we propose a robust loss function specifically designed for high-dimensional MVG outputs. The proposed MVG-CRPS loss function has a closed-form expression based on the neural network outputs, making it easily integrable into deep learning models. We evaluate MVG-CRPS on two probabilistic forecasting tasks—multivariate autoregressive and univariate sequence-to-sequence (Seq2Seq) forecasting—both involving observations following MVG distribution. Experimental results on real-world datasets demonstrate that MVG-CRPS achieves both robustness and efficiency, offering enhanced accuracy and uncertainty quantification in probabilistic forecasting.
@@ 1 INTRODUCTION
Probabilistic forecasting aims to capture the uncertainty inherent in time series data, offering a distribution of possible future values rather than single-point estimates provided by deterministic forecasts. This approach is invaluable in various domains such as finance (Groen et al.,, 2013), weather prediction (Palmer,, 2012), and health care management (Jones and Spiegelhalter,, 2012), where understanding the range of potential outcomes is crucial for risk assessment and informed decision-making.
@@
To achieve accurate probabilistic forecasts, it is crucial to develop a well-specified forecasting model and employ an effective evaluation metric during training. While the development of probabilistic forecasting models has received significant attention in recent years, the advancement of evaluation metrics has been less explored (Bjerregård et al.,, 2021). A common approach for evaluating a forecaster involves applying scoring rules that assign a numerical score to each forecast, which are then collected across a test dataset. The choice of scoring rule typically depends on the type of forecasting task—whether it is deterministic or probabilistic, univariate or multivariate—and the nature of the variable being forecasted, such as categorical or continuous. In this paper, we focus on continuous variables within the probabilistic forecasting setting.
@@
Although scoring rules for univariate probabilistic forecasting are well-established, their multivariate counterparts remain less developed (Panagiotelis et al.,, 2023). Multivariate time series data, involving multiple interdependent variables, present additional challenges such as increased complexity, higher dimensionality, and intricate cross-variable interactions. Traditional evaluation metrics like the CRPS (Matheson and Winkler,, 1976) and the log-score have been effective in univariate contexts but often fall short in multivariate scenarios due to issues like sensitivity to outliers and computational inefficiency. For example, when extended to the multivariate case, the CRPS lacks a closed-form expression, leading to computational inefficiency; additionally, the log-score heavily penalizes the tails of the data distribution, making it more sensitive to outliers (Gebetsberger et al.,, 2018; Bjerregård et al.,, 2021).
@@
To address these challenges, we propose a robust loss function specifically designed for multivariate probabilistic time series forecasting with a Gaussian distribution output layer. Our method utilizes a PCA whitening transformation to decorrelate the multivariate time series variables into a new random vector with zero mean and identity covariance matrix. Consequently, each component of this vector follows a standard Gaussian distribution, allowing us to apply the closed-form expression of the CRPS for evaluation. This approach not only overcomes the lack of a closed-form CRPS in the multivariate case but also reduces sensitivity to outliers and extreme tails, enhancing both robustness and computational efficiency. The key contributions of our work are:
@@
  • We introduce a loss function that is less sensitive to outliers and extreme tails of the data distribution. This robustness ensures that the forecaster focuses on the overall data structure rather than being unduly influenced by anomalies.
@@
  • The proposed loss function has a closed-form expression, allowing for the analytical computation of derivatives. This feature is particularly advantageous for training deep learning models, as it enables seamless integration with backpropagation algorithms and reduces computational overhead.
@@
  • We conduct extensive experiments using deep probabilistic forecasting models on real-world datasets. Our results show that the proposed loss function outperforms traditional scoring rules in terms of both accuracy and robustness.
@@ 2.1 Probabilistic Forecasting
Probabilistic forecasting concentrates on modeling the probability distribution of target variables, as opposed to deterministic forecasting, which offers only single-point estimates. This approach is vital for capturing the uncertainty inherent in time series data, enabling better risk assessment and decision-making. There are two primary methodologies for probabilistic forecasting: parametric (e.g., through probability density functions (PDFs)) and non-parametric (e.g., through quantile functions) (Benidis et al.,, 2022).
PDF-based methods involve assuming a specific probability distribution (e.g., Gaussian, Poisson) and using neural networks to estimate the parameters of that distribution. For example, DeepAR (Salinas et al.,, 2020) utilizes a recurrent neural network (RNN) to model hidden state transitions and predict distribution parameters at each time step. Its multivariate extension, GPVar (Salinas et al.,, 2019), employs a Gaussian copula to convert observations into Gaussian variables, thereby assuming a joint MVG distribution. This technique captures dependencies among multiple time series, which is crucial for accurate multivariate probabilistic forecasting. In the Gaussian case, one can further introduce a batch dimension and use Generalized Least Square (GLS) method to capture the autocorrelation and cross-correlation among different time steps (Zheng et al.,, 2024; Zheng and Sun,, 2024).
Neural networks can also generate parameters for more complex probabilistic models. The deep state space model (Rangapuram et al.,, 2018) employs an RNN to produce parameters for state space models, effectively modeling both temporal dynamics and forecast uncertainty. The normalizing Kalman filter (de Bézenac et al.,, 2020) integrates normalizing flows with linear Gaussian state space models to handle nonlinear dynamics and compute the PDFs of observations. In this framework, RNNs generate state space model parameters at each time step, and normalizing flows transform the output into observations, allowing for more flexible distribution modeling. Normalizing flows (Rasul et al., 2021b, ) model complex, non-Gaussian distributions by transforming a simple base distribution through a sequence of invertible functions. Diffusion models (Rasul et al., 2021a, ) offer another strategy by defining a stochastic process that incrementally adds noise to the data and learning to reverse this process to generate samples from the desired distribution. Copula-based methods have also been explored to model dependencies between multiple time series. Studies by Drouin et al., (2022) and Ashok et al., (2024) use copulas to construct multivariate distributions by specifying individual marginal distributions and a copula function that captures the dependence structure. This allows for flexible modeling of inter-variable relationships, which is essential in multivariate probabilistic forecasting.
@@
Quantile-based approaches focus on predicting specific quantiles of the target distribution without relying on strict parametric assumptions. For instance, MQ-RNN (Wen et al.,, 2017) produces quantile forecasts using a Seq2Seq RNN architecture. By forecasting multiple quantiles, these models can reconstruct the entire target distribution, which is particularly useful for capturing asymmetric or heavy-tailed behaviors.
Among the existing works that model the full probability distribution using a parametric approach, most employ the log-score as the loss function for optimizing model parameters.
@@ 2.2 Scoring Rules
Scoring rules are fundamental tools used to evaluate the quality of probabilistic forecasts by assigning numerical scores based on the predicted probability distributions and the observed outcomes. They not only assess the accuracy and calibration of forecasts but can also be utilized in the estimation of model parameters during training. A scoring rule is considered proper if the expected score is minimized when the forecasted probability distribution $p$ matches the true distribution $q$ of the observations. Formally, a scoring rule $s(p,q)$ is proper if the divergence $d(p,q)=s(p,q)-s(q,q)$ is non-negative and it is strictly proper if $d(p,q)=0$ implies $p=q$ (Bröcker,, 2009).
@@
One of the simplest and most widely used scoring rules in both univariate and multivariate settings is the log-score, also known as the negative log-likelihood (NLL). It is defined as the negative logarithm of the predictive density evaluated at the observed value. This score is particularly common when a parametric form of the predictive density is available (Panagiotelis et al.,, 2023). The log-score is a strictly proper scoring rule and has several desirable properties, such as consistency and sensitivity to the entire distribution. However, it is known to lack robustness because it heavily penalizes forecasts that assign low probability to the observed outcome, making it highly sensitive to outliers and extreme events (Gneiting et al.,, 2007).
@@
To address the sensitivity of the log-score to outliers, CRPS has been proposed as a more robust alternative, especially in univariate settings. CRPS measures the difference between the cumulative distribution functions (CDFs) of the forecast and the observation, effectively integrating the absolute error over all possible threshold values. It can be interpreted as a generalized version of the mean absolute error (MAE) (Gneiting et al.,, 2005). A key distinction between the log-score and CRPS is that CRPS grows linearly with the normalized prediction error, whereas the log-score grows quadratically. Consequently, the log-score assigns harsher penalties to poor probabilistic forecasts, increasing its sensitivity to outliers.
@@
CRPS has been effectively used for parameter estimation in statistical models. For example, the method of minimum CRPS estimation has been introduced for fitting ensemble model output statistics (EMOS) coefficients, optimizing the forecasts by directly minimizing the CRPS rather than maximizing the likelihood (Gneiting et al.,, 2005). This approach can lead to better-calibrated predictive distributions that are not excessively wide or overdispersed.
Extending scoring rules to multivariate probabilistic forecasting introduces additional complexity due to the dependencies among variables and higher dimensionality. While the log-score remains applicable in multivariate contexts when a parametric form of the joint predictive density is available, it continues to suffer from sensitivity to outliers.
@@
The Energy Score (Gneiting and Raftery,, 2007) is a commonly used multivariate generalization of the CRPS. It assesses the discrepancy between the forecasted and observed distributions by considering the expectations of distances between random vectors. It is reported that the energy score is effective at detecting errors in the mean of the forecast distribution but is less sensitive to errors in the variance and, more critically, to misspecifications in the correlation structure among variables (Pinson and Tastu,, 2013; Alexander et al.,, 2024). The absence of a closed form expression also necessitates the use of Monte Carlo simulations to approximate the score by drawing samples from the predictive distribution, which can be computationally intensive (Panagiotelis et al.,, 2023).
@@
Other multivariate scoring rules include the Variogram Score (Scheuerer and Hamill,, 2015) and the Dawid-Sebastiani Score (Wilks,, 2020). The variogram score focuses on assessing the spatial or temporal dependence structure by comparing differences between pairs of observations and forecasts. The Dawid-Sebastiani score combines aspects of both the log-score and variance to evaluate forecasts. However, like the energy score, these metrics can face computational challenges and may require approximations when closed-form solutions are unavailable. We refer readers to the comprehensive reviews by Gneiting and Katzfuss, (2014) and Tyralis and Papacharalampous, (2024) for further details.
@@ 3.1 Multivariate Probabilistic Forecasting
Probabilistic forecasting aims to estimate the joint distribution over a collection of future quantities based on a given history of observations (Gneiting and Katzfuss,, 2014). Denote the time series vector at a time point $t$ as $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$, where $N$ is the number of series. The problem of probabilistic forecasting can be formulated as $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$, $P$ is the conditioning range, $Q$ is the prediction range, and $T$ is the time point that splits the conditioning range and prediction range. $\mathbf{x}_{t}$ are some known covariates for both past and future time steps.
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ are the functions mapping $\mathbf{h}_{t}$ to the mean and covariance parameters. The log-likelihood of the distribution given observed time series data up to time point $T$ can be used as the loss function for optimizing a DL model
@@
where $F$ is the CDF of the predicted variable, $z$ is the observation, and $Z$ and $Z^{\prime}$ are independent random variables both associated with the CDF $F$. The CRPS has a closed-form expression when evaluating a Gaussian-distributed variable $z\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$ (Gneiting et al.,, 2005):
$\operatorname{CRPS}\left(F,z\right)=z\left(2F\left(z\right)-1\right)+2f\left(z\right)-\frac{1}{\sqrt{\pi}},$ (8)
$\operatorname{CRPS}\left(F_{\mu,\sigma},z\right)=\sigma\operatorname{CRPS}\left(F,\frac{z-\mu}{\sigma}\right),$ (9)
where $F_{\mu,\sigma}\left(z\right)=F\left(\frac{z-\mu}{\sigma}\right)$, $F$ and $f$ are the CDF and PDF of the standard Gaussian distribution.
The CRPS has been shown to be a more robust alternative to the log-score as a loss function in univariate problems, particularly because the negative log-likelihood in Eq. (6) is quadratic in form (Gneiting et al.,, 2005; Rasp and Lerch,, 2018; Murad et al.,, 2021; Chen et al.,, 2024). Moreover, the CRPS can directly replace the log-score, providing analytical gradients with respect to $\mu$ and $\sigma$ for backpropagation. However, for a MVG distribution, the CRPS does not have a widely used closed-form expression.

### v2

@@ Abstract
Multivariate Gaussian (MVG) distributions are central to modeling correlated continuous variables in probabilistic forecasting. Neural forecasting models typically parameterize the mean vector and covariance matrix of the distribution using neural networks, optimizing with the log-score (negative log-likelihood) as the loss function. However, the sensitivity of the log-score to outliers can lead to significant errors in the presence of anomalies. Drawing on the continuous ranked probability score (CRPS) for univariate distributions, we propose MVG-CRPS, a strictly proper scoring rule for MVG distributions. MVG‐CRPS admits a closed‐form expression in terms of neural network outputs, thereby integrating seamlessly into deep learning frameworks. Experiments on real-world datasets across multivariate autoregressive and univariate sequence-to-sequence (Seq2Seq) forecasting tasks show that MVG-CRPS improves robustness, accuracy, and uncertainty quantification in probabilistic forecasting.
@@ 1 Introduction
Probabilistic forecasting aims to capture the uncertainty inherent in time series data, offering a distribution of possible future values rather than single-point estimates provided by deterministic forecasts. This approach is invaluable in various domains such as finance (Groen et al., 2013), weather prediction (Palmer, 2012), and health care management (Jones & Spiegelhalter, 2012), where understanding the range of potential outcomes is crucial for risk assessment and informed decision-making.
@@
To achieve accurate probabilistic forecasts, it is crucial to develop a well-specified forecasting model and employ an effective evaluation metric during training. While the development of probabilistic forecasting models has received significant attention in recent years, the advancement of evaluation metrics has been less explored (Bjerregård et al., 2021). A common approach for evaluating a forecaster involves applying scoring rules that assign a numerical score to each forecast, which are then collected across a test dataset. The choice of scoring rule typically depends on the type of forecasting task—whether it is deterministic or probabilistic, univariate or multivariate—and the nature of the variable being forecasted, such as categorical or continuous. In this paper, we focus on continuous variables within the probabilistic forecasting setting.
Figure 1: A motivating example illustrating how MVG-CRPS improves predictive performance by limiting outlier influence and reducing training time through its sampling-free design. The energy score is computed using sample sizes of 50, 100, and 150.
@@
Although scoring rules for univariate probabilistic forecasting are well-established, their multivariate counterparts remain less developed (Panagiotelis et al., 2023). Multivariate time series data, characterized by multiple interconnected variables, present additional challenges such as increased complexity, higher dimensionality, and intricate cross-variable interactions. Traditional evaluation metrics like the CRPS (Matheson & Winkler, 1976) and the log-score have been effective in univariate contexts but often fall short in multivariate scenarios due to issues like sensitivity to outliers and computational inefficiency. For example, when extended to the multivariate case, the CRPS lacks a closed-form expression, leading to computational inefficiency; additionally, the log-score heavily penalizes the tails of the data distribution, making it more sensitive to outliers (Gebetsberger et al., 2018; Bjerregård et al., 2021).
@@
To overcome these challenges, we introduce MVG-CRPS, a strictly proper scoring rule utilized as a loss function for multivariate probabilistic forecasting with a Gaussian distribution output layer. Our method utilizes a PCA whitening transformation to decorrelate the multivariate variables into a new random vector with zero mean and identity covariance matrix. Consequently, each component of this vector follows a standard Gaussian distribution, allowing us to apply the closed-form expression of the CRPS for evaluation. This approach not only overcomes the lack of a closed-form CRPS in the multivariate case but also reduces sensitivity to outliers and extreme tails, enhancing both robustness and computational efficiency. The advantages of our approach are illustrated in Fig. 1, where the model trained with MVG-CRPS achieves higher accuracy while significantly reducing training time. The key contributions of our work are:
@@
  • We propose a scoring rule for multivariate probabilistic forecasting that is less sensitive to outliers and extreme tails of the data distribution. Under the Gaussian assumption, we prove that this scoring rule is strictly proper. Its robustness ensures that the resulting forecaster prioritizes the overall data structure rather than being disproportionately affected by anomalies.
@@
  • The proposed scoring rule has a closed-form expression, allowing for the analytical computation of derivatives. This feature is particularly advantageous for training deep learning models, as it enables seamless integration with backpropagation algorithms and reduces computational overhead.
@@
  • We perform extensive experiments with deep probabilistic forecasting models on real-world datasets. Our results demonstrate that the proposed method balances accuracy and efficiency more effectively than common scoring rules.
@@ 2.1 Probabilistic Forecasting
Probabilistic forecasting concentrates on modeling the probability distribution of target variables, as opposed to deterministic forecasting, which offers only single-point estimates. This approach is vital for capturing the uncertainty inherent in time series data, enabling better risk assessment and decision-making. There are two primary distributional regression methodologies for probabilistic forecasting: parametric (e.g., through probability density functions (PDFs)) and non-parametric (e.g., through quantile functions) (Benidis et al., 2022).
@@
Non-parametric approaches typically predict specific quantiles of the target distribution, avoiding the need for strict parametric assumptions. For example, MQ-RNN (Wen et al., 2017) generates quantile forecasts using a Seq2Seq RNN architecture. By forecasting multiple quantiles, these models provide an approximation of the target distribution, making them especially effective for capturing asymmetries or heavy-tailed behaviors.
PDF-based methods involve assuming a specific probability distribution (e.g., Gaussian, Poisson) and using neural networks to estimate the parameters of that distribution. For example, DeepAR (Salinas et al., 2020) utilizes a recurrent neural network (RNN) to model hidden state transitions and predict Gaussian distribution parameters at each time step. Its multivariate extension, GPVar (Salinas et al., 2019), employs a Gaussian copula to convert observations into Gaussian variables, thereby assuming a joint MVG distribution. This technique captures dependencies among multiple time series, which is crucial for accurate multivariate probabilistic forecasting. In the Gaussian case, one can further introduce a batch dimension and use generalized least squares (GLS) method to capture the autocorrelation and cross-correlation among different time steps (Zheng et al., 2024; Zheng & Sun, 2024).
Neural networks can also parameterize more complex probabilistic models, enabling greater flexibility and expressiveness. Examples include state space models (SSMs) (Rangapuram et al., 2018; de Bézenac et al., 2020), normalizing flows (NFs) (Rasul et al., 2021b), and diffusion models (Rasul et al., 2021a). Copula-based methods have also been explored to model dependencies between multiple time series. Studies by Drouin et al. (2022) and Ashok et al. (2024) use copulas to construct multivariate distributions by specifying individual marginal distributions and a copula function that captures the dependence structure. This allows for flexible modeling of inter-variable relationships, which is essential in multivariate probabilistic forecasting. Among existing approaches that model the full probability distribution, the majority use the log-score as the loss function to optimize model parameters.
@@ 2.2 Scoring Rules
Scoring rules are fundamental tools used to evaluate the quality of probabilistic forecasts by assigning numerical scores based on the predicted probability distributions and the observed outcomes. A scoring rule is considered proper if the expected score is minimized when the predicted probability distribution $p$ matches the true distribution $q$ of the observations. Formally, a scoring rule $s(p,q)$ is proper if the divergence $d(p,q)=s(p,q)-s(q,q)$ is non-negative and it is strictly proper if $d(p,q)=0$ implies $p=q$ (Bröcker, 2009).
@@
One of the most widely used scoring rules in both univariate and multivariate settings is the log-score, also known as the negative log-likelihood (NLL). It is defined as the negative logarithm of the predictive density evaluated at the observed value. This score is particularly common when a parametric form of the predictive density is available (Panagiotelis et al., 2023). The log-score is a strictly proper scoring rule and has several desirable properties, such as consistency and sensitivity to the entire distribution. However, it is known to lack robustness because it heavily penalizes forecasts that assign low probability to the observed outcome, making it highly sensitive to outliers and extreme events (Gneiting et al., 2007).
@@
To address the sensitivity of the log-score to outliers, CRPS has been proposed as a more robust alternative, especially in univariate settings (Rasp & Lerch, 2018). CRPS measures the difference between the predictive CDF (cumulative distribution function) and the observation, effectively integrating the absolute error over all possible threshold values. It can be interpreted as a generalized version of the mean absolute error (MAE) (Gneiting et al., 2005). A key distinction between the log-score and CRPS is that CRPS grows linearly with the prediction error, while the log-score penalizes errors exponentially as the prediction deviates further from the observed value. Consequently, the log-score assigns harsher penalties to poor probabilistic forecasts, increasing its sensitivity to outliers.
@@
CRPS has been effectively used for parameter estimation in probabilistic models (Gneiting et al., 2005; Olivares et al., 2023; Lang et al., 2024). For example, the method of minimum CRPS estimation has been introduced for fitting ensemble model output statistics (EMOS) coefficients, optimizing the forecasts by directly minimizing the CRPS rather than maximizing the likelihood (Gneiting et al., 2005). This approach can lead to better-calibrated predictive distributions that are not excessively wide or overdispersed.
@@
Extending scoring rules to multivariate forecasting introduces additional challenges due to variable dependencies and increased dimensionality. Although the log-score can be applied in multivariate settings, its sensitivity to outliers persists. The energy score (ES) (Gneiting & Raftery, 2007), a widely used multivariate extension of the CRPS, evaluates the discrepancy between predictive and observed distributions by measuring the expected distances between random vectors. While ES effectively detects errors in the forecast mean, it is less sensitive to variance errors and, more critically, to misspecifications in the correlation structure among variables (Pinson & Tastu, 2013; Alexander et al., 2024). The absence of a closed form expression also necessitates the use of Monte Carlo simulations to approximate the score by drawing samples from the predictive distribution, which can be computationally intensive (Panagiotelis et al., 2023).
@@
Other multivariate scoring rules include the Variogram Score (Scheuerer & Hamill, 2015) and the Dawid-Sebastiani Score (Wilks, 2020). These metrics also face computational challenges and may require approximations when closed-form solutions are unavailable. We refer readers to the comprehensive reviews by Gneiting & Katzfuss (2014), Ziel & Berk (2019), and Tyralis & Papacharalampous (2024) for further details.
@@ 3.1 Multivariate Probabilistic Forecasting
Probabilistic forecasting aims to estimate the joint distribution over a collection of future quantities based on a given history of observations (Gneiting & Katzfuss, 2014). Denote the time series vector at a time point $t$ as $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$, where $N$ is the number of series. The problem of probabilistic forecasting can be formulated as $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$, $P$ is the conditioning range, $Q$ is the prediction range, and $T$ is the time point that splits the conditioning range and prediction range. $\mathbf{x}_{t}$ are some known covariates for both past and future time steps.
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ are the functions mapping $\mathbf{h}_{t}$ to the mean and covariance parameters. The log-likelihood of the distribution given observed time series data up to time point $T$ can be used as the loss function for optimizing a DL model:
@@
where $F$ is the predictive CDF, $z$ is the observation, and $Z$ and $Z^{\prime}$ are independent random variables both associated with the CDF $F$. The CRPS has a closed-form expression when evaluating a Gaussian-distributed variable $z\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$ (Gneiting et al., 2005):
$\operatorname{CRPS}\left(\Phi,z\right)=z\left(2\Phi\left(z\right)-1\right)+2\varphi\left(z\right)-\frac{1}{\sqrt{\pi}},$ (8)
$\operatorname{CRPS}\left(F_{\mu,\sigma},z\right)=\sigma\operatorname{CRPS}\left(\Phi,\frac{z-\mu}{\sigma}\right),$ (9)
where $F_{\mu,\sigma}\left(z\right)=\Phi\left(\frac{z-\mu}{\sigma}\right)$, $\Phi$ and $\varphi$ are the CDF and PDF of the standard Gaussian distribution.
The CRPS has been shown to be a more robust alternative to the log-score as a loss function in many problems (Gneiting et al., 2005; Rasp & Lerch, 2018; Murad et al., 2021; Chen et al., 2024). We observe that the log-score can grow arbitrarily large in magnitude when a single outlier disproportionately influences the loss function, owing to the unbounded nature of the logarithmic function (Eq. (4) and Eq. (6)). Additionally, the quadratic form of the error terms in the Gaussian likelihood also makes it sensitive to outliers (e.g., $\epsilon_{i,t}^{2}$ in Eq. (6)). In contrast, CRPS evaluates the entire predictive distribution rather than concentrating solely on the likelihood of individual data points (Eq. (8)). Moreover, the CRPS can directly replace the log-score, providing analytical gradients with respect to $\mu$ and $\sigma$ for backpropagation. However, for a MVG distribution, the CRPS does not have a widely used closed-form expression.